OpenStudy (anonymous):

Part 1: Using complete sentences, explain how to find the equation of the line, in standard form and slope–intercept form, passing through (–1, –7) and (1, –1). Part 2: Compare the benefits of writing an equation in standard form to the benefits of writing an equation in slope–intercept form.

OpenStudy (abb0t):

\[slope = m = \frac{ y_2 - y_1 }{ x_2-x_1 }\] line: y = mx + b where b = y-coordinate Also, \[(y-y_0)=m(x-x_0)\]